# Guided simulation of conditioned chemical reaction networks

### v1

@@ 2.1. Distribution of reaction times
With each reaction $\ell\in{\mathcal{R}}$, we associate a reaction time $\tau_{\ell}$. If $\lambda_{\ell}$ is constant in time, it follows from (2.1) that $\tau_{\ell}\mid X(s)=x\sim\mathrm{Exp}\left(\lambda_{\ell}(x)\right)$. This implies that the time $\tau=\min_{\ell\in{\mathcal{R}}}\tau_{\ell}$ that the first reaction occurs satisfies $\tau\mid X(s)=x\sim\mathrm{Exp}\left(\sum_{\ell\in{\mathcal{R}}}\lambda_{\ell}(x)\right)$. If the reaction times are inhomogeneous in time and 2.2 applies, the reaction times satisfy, see e.g. Gillespie, (1992),
@@
Similarly the time for the first reaction to occur has distribution (2.4) with intensity function $\sum_{\ell\in{\mathcal{R}}}\lambda_{\ell}$. Moreover, the probability distribution of the first reaction in any subset $R\subseteq{\mathcal{R}}$ also satisfies (2.4) with intensity function $\sum_{\ell\in R}\lambda_{\ell}$.
@@ 2.2. Chemical master equation and chemical Langevin equation
with initial condition $p(s,x;s,y)=1\{y=x\}$. It is well-known, see e.g. Li, (2020) that solutions to the chemical master equation can be approximated by the transition density of solutions to the Chemical Langevin Equation (CLE), which is the SDE
@@
where (assuming the reactions to be numbered $1,\ldots,p$)
$\begin{split}b_{\mathrm{CLE}}(t,x)&=\sum_{\ell=1}^{p}\lambda_{\ell}(t,x)\xi_{\ell}\\ \sigma_{\mathrm{CLE}}(t,x)&=\begin{bmatrix}\xi_{1}&\ldots&\xi_{p}\end{bmatrix}\sqrt{\mbox{diag}(\lambda_{1}(t,x),\ldots,\lambda_{p}(t,x))}\end{split}$ (2.6)
and $W$ is an independent ${\mathbb{R}}^{d}$-valued Brownian motion.
@@ 2.3. Examples
Our simplest example models a population of initial size $x_{0}$ in which an individual dies in a time interval $(t,t+\Delta)$ with probability $c\Delta+o(\Delta)$ for some constant $c>0$ and $\Delta$ small. Such a process is modelled as chemical reaction process with just one specie and one reaction, namely $(t,x)\mapsto cx$ with $\xi=-1$.
@@
Here $(x_{1},x_{2},x_{3})$ refer to species counts of $(S,E,SE)$. Interesting aspects of this example are firstly that the fourth component ($P$) only appears in reaction ($3$) where $1$ is added and therefore is monotonically increasing and secondly that there are absorbig states such as $x=(0,x_{2},0,x_{4})$ where the process is killed as all reaction rates are zero. A realization of this process can be found in Figure 2.
Figure 2. Realization of the forward process of 2.5 using $\kappa_{1}=5$, $\kappa_{2}=5$, $\kappa_{3}=3$ and initial position $x_{0}=(12,10,10,10)$.
3. Doob $h$-transform and guided Markov processes
We assume $x_{0}$ to be known and that $v=LX(T)$ is observed, where $L\in{\mathbb{R}}^{m\times d}$ is of full row rank and $m\leq d$. In case $m=d$, we assume without loss of generality that $L=I$. The extension to multiple observations will follow in a straightforward way from this case. We aim to sample $X$ on $[0,T]$, conditional on the event $\{LX(T)=v\}$. Our approach crucially depends on results from Palmowski & Rolski, (2002) on exponential change of measure.
For a suitable $f\colon[0,T)\times{\mathbb{S}}\to{\mathbb{R}}$ define
$D_{t}^{f}=\frac{f(t,X(t))}{f(0,x_{0})}\exp\left(-\int_{0}^{t}\frac{{\mathcal{A}}f}{f}(s,X(s))\,\mathrm{d}s\right),\qquad t\in[0,T).$ (3.1)
@@ 3. Doob $h$-transform and guided Markov processes
We write $g\in{\mathcal{G}}$ if $g\in{\mathcal{D}}({\mathcal{A}})$ is a strictly positive function such that $D_{t}^{g},\,t\in[0,T)$ is a martingale adapted to $({\mathcal{F}}_{t})_{t}$.
Let ${\mathbb{P}}_{t}$ denote the law of the process $X$, restricted to ${\mathcal{F}}_{t}$, let $g\in{\mathcal{G}}$ and define the change of measure
$\,\mathrm{d}{\mathbb{P}}_{t}^{g}=D_{t}^{g}\,\mathrm{d}{\mathbb{P}}_{t},\qquad t\in[0,T).$ (3.2)
By formula (1.2) of Palmowski & Rolski, (2002), the process under ${\mathbb{P}}_{t}^{g}$ has generator
${\mathcal{L}}_{t}^{g}f(x)=\frac{1}{g(t,x)}\left[{\mathcal{L}}_{t}\left(fg\right)(t,x)-f(x){\mathcal{L}}_{t}g(t,x)\right]=\sum_{\ell}\lambda_{\ell}^{g}(t,x)\left(f(x+\xi_{\ell})-f(x)\right),$ (3.3)
with
$\lambda_{\ell}^{g}(t,x)=\alpha^{g}_{\ell}(t,x)\lambda_{\ell}(t,x)$ (3.4)
and
$\alpha^{g}_{\ell}(t,x)=\frac{g(t,x+\xi_{\ell})}{g(t,x)},\qquad(\ell,t,x)\in{\mathcal{R}}\times[0,T)\times{\mathbb{S}}.$ (3.5)
@@
Some authors refer to this process as the twisted process (see e.g. Del Moral & Penev, (2017)). Under ${\mathbb{P}}_{t}^{g}$, the process is a chemical reaction process with the same jumps as under ${\mathbb{P}}_{t}$ but with adjusted jump intensities in (3.4). The jump intensities in (3.4) suggest that the process obtained from the change of measure in (3.1) is invariant under multiplying $g$ by a function of time. This is formalized in 3.3.
@@
Suppose $g\in{\mathcal{G}}$ and let $c\colon[0,T]\to{\mathbb{R}}_{+}$ be a differentiable function. Then $D_{t}^{cg}=D_{t}^{g}$.
@@
Upon denoting the transition probabilities of $X$ by $p$, i.e. ${\mathbb{P}}(X_{T}\in A\mid X_{t}=x)=\sum_{y\in A}p(t,x;T,y)$ for $A\subseteq{\mathbb{S}}$, it follows from Example 2.4 of Corstanje et al., (2023) that the conditioned process $\left(X\mid LX(T)=v\right)$ is obtained by taking $g=h$ with
$h(t,x)=\sum_{\zeta\in L^{-1}v}p(t,x;T,\zeta),$ (3.6)
where $L^{-1}v=\{y\in{\mathbb{S}}:Ly=v\}$. Throughout, we assume $h(0,x_{0})>0$. Note that $h$ satisfies Kolmogorov’s backward equation and therefore $h\in{\mathcal{G}}$ and ${\mathcal{A}}h=0$. Hence,
$\,\mathrm{d}{\mathbb{P}}_{t}^{h}=\frac{h(t,X(t))}{h(0,x_{0})}\,\mathrm{d}{\mathbb{P}}_{t},\qquad t<T.$
Intuitively, ${\mathbb{P}}_{t}^{h}$ gives more mass to paths where $h(t,X(t))$ is large. Absolute continuity of ${\mathbb{P}}_{t}^{h}$ with respect to ${\mathbb{P}}_{t}$ won’t hold under the limit $t\uparrow T$ as under ${\mathbb{P}}^{h}_{t}$ the process is conditioned on the event $\{LX_{T}=v\}$. Unfortunately, $h$ is intractable as the transition probabilities $p$ are only known in closed-form in very specific cases. Therefore, consider $g\in{\mathcal{G}}$ that acts as a tractable substitute for $h$. Then, as a consequence, the process $X$ can be simulated under the measure ${\mathbb{P}}_{t}^{g}$ and moreover, for $t<T$,
$\frac{\,\mathrm{d}{\mathbb{P}}^{h}_{t}}{\,\mathrm{d}{\mathbb{P}}^{g}_{t}}(X)=\frac{D_{t}^{h}}{D_{t}^{g}}=\frac{h(t,X(t))}{g(t,X(t))}\frac{g(0,x_{0})}{h(0,x_{0})}\Psi_{t}^{g}(X),$ (3.7)
where
$\Psi_{t}(X)=\exp\left(\int_{0}^{t}\frac{{\mathcal{A}}g}{g}(s,X(s))\,\mathrm{d}s\right).$ (3.8)
To evaluate $\Psi$, we note that a direct computation yields
Theorem 3.4 (Limiting case of (3.7) as $t\uparrow T$).
Suppose $h$ is as in (3.6) and $g\in{\mathcal{G}}$ is such both $\frac{\partial}{\partial t}\log g$ and $\sum_{\ell\in{\mathcal{R}}}\lambda_{\ell}^{g}$ are bounded. Then ${\mathbb{P}}_{T}^{h}\ll{\mathbb{P}}_{T}^{g}$ with
$\frac{\,\mathrm{d}{\mathbb{P}}_{T}^{h}}{\,\mathrm{d}{\mathbb{P}}_{T}^{g}}(X)=\frac{g(0,x_{0})}{h(0,x_{0})}\frac{\Psi_{T}(X)}{g(T,X(T-))}\mathbf{1}_{\{LX(T)=v\}},$ (3.10)
where $X(T-)=\lim_{t\uparrow T}X(t)$ $\mathbb{P}^{g}$-almost surely.
Proof.
First note that, as $\sum_{\ell\in{\mathcal{R}}}\lambda_{\ell}^{g}$ is bounded. $X$ has finitely many jumps in $[0,T]$ and $X(T-)$ and $X(T)$ exist ${\mathbb{P}}^{g}$-almost surely. By (3.7), ${\mathbb{P}}_{t}^{h}\ll{\mathbb{P}}_{t}^{g}$ for $t<T$ and hence for any ${\mathcal{F}}_{s}$-measurable function $f_{s}$ for $s<T$ and $t\in(s,T)$,
${\mathbb{E}}^{h}_{t}f_{s}(X)={\mathbb{E}}_{t}^{g}\left[f_{s}(X)\frac{g(0,x_{0})}{h(0,x_{0})}\frac{h(t,X(t))}{g(t,X(t))}\exp\left(\int_{0}^{t}\frac{{\mathcal{A}}g}{g}(s,X(s))\,\mathrm{d}s\right)\right].$ (3.11)
Now note that as $t\uparrow T$, ${h}(t,X(t))={\mathbb{E}}\left[\mathbf{1}\{LX(T)=v\}\mid{\mathcal{F}}_{t}\right]\to{\mathbb{E}}\left[\mathbf{1}\{LX(T)=v\}\mid{\mathcal{F}}_{T}\right]=\mathbf{1}\{LX(T)=v\}$. Hence, taking $\lim_{t\uparrow T}$ on both sides of (3.11) and interchanging limit and expectation finishes the proof. To see that the swap is allowed, note that it follows from (3.9) that $\frac{{\mathcal{A}}g}{g}$ is bounded. Moreover, as $g$ is differentiable in $t$ (as $g\in{\mathcal{G}}$), the maps $t\mapsto g(t,x)$ are continuous for all $x\in{\mathbb{S}}$ and thus $t\mapsto g(t,X(t))$ is ${\mathbb{P}}^{g}$-almost surely cadlag on $[0,T]$ with finitely many jumps and we can find a positive lower bound for it. ∎
Corollary 3.5.
The likelihood of $x_{0}$ based on the observation $v$ is given by
$h(0,x_{0})=g(0,x_{0}){\mathbb{E}}_{T}^{g}\left[\frac{\Psi_{T}(X)}{g(T,X(T-)}\mathbf{1}_{\{LX(T)=v\}}\right].$ (3.12)
Proof.
We present one class of functions that satisfies the conditions of 3.4
Proposition 3.6.
Suppose $g\in{\mathcal{D}}({\mathcal{A}})$ is such that for all $x\in{\mathbb{S}}$ the map $t\mapsto g(t,x)$ is continuously differentiable on $[0,T]$ and there exist $\delta,K$ such that $0<\delta\leq g(t,x)\leq K<\infty$ for all $(t,x)\in[0,T]\times{\mathbb{S}}$. Then, by 3.4, ${\mathbb{P}}_{T}^{h}\ll{\mathbb{P}}_{T}^{g}$.
Proof.
As $g$ is bounded away from zero, Lemma 3.2 of Chapter 4 in Ethier & Kurtz, (2009) implies that $D_{t}^{g}$ is a martingale. Hence $g\in{\mathcal{G}}$. Moreover, $\partial_{t}\log g=g^{-1}\partial_{t}g$ is bounded as $g\geq\delta$ and $\partial_{t}g(t,x)$ is continuous on $[0,T]$ for fixed $x$. Lastly, for all $(t,x)$ we have that $\sum_{\ell\in{\mathcal{R}}}\lambda_{\ell}^{g}(t,x)\leq(K/\varepsilon)\sum_{\ell\in{\mathcal{R}}}\lambda_{\ell}(t,x)$, which is bounded by Assumption (2.1c). ∎
The following proposition is useful to numerically evaluate the likelihood of a sampled guided process.
@@
$\begin{split}&\frac{g(0,x_{0})}{g(T,x(T-))}\exp\left(\int_{0}^{T}\frac{{\mathcal{A}}g}{g}(s,x(s))\,\mathrm{d}s\right)\\ &\qquad=\left(\prod_{j=0}^{N-2}\frac{1}{\alpha_{j}(t_{j+1})}\right)\exp\left(\sum_{j=0}^{N-1}\int_{t_{j}}^{t_{j+1}}\sum_{\ell\in{\mathcal{R}}}\lambda_{\ell}(s,x_{j})\left[\alpha_{j}(s)-1\right]\,\mathrm{d}s\right)\end{split}$
@@
We have
$\exp\left(\int_{0}^{T}\partial_{s}\log g(s,x(s))\,\mathrm{d}s\right)$ $=\exp\left(\sum_{j=0}^{N-1}\int_{t_{j}}^{t_{j+1}}\frac{\,\mathrm{d}}{\,\mathrm{d}s}\log g(s,x_{j})\,\mathrm{d}s\right)$
$=\frac{g(t_{N},x_{N-1})}{g(t_{0},x_{0})}\prod_{j=0}^{N-2}\frac{g(t_{j+1},x_{j})}{g(t_{j+1},x_{j+1})}.$
Moreover,
$\int_{0}^{T}\frac{{\mathcal{L}}g}{g}(s,x(s))\,\mathrm{d}s=\sum_{j=0}^{N-1}\int_{t_{j}}^{t_{j+1}}\sum_{\ell\in{\mathcal{R}}}\lambda_{\ell}(s,x_{j})\left[\frac{g(s,x_{j+1})}{g(s,x_{j})}-1\right]\,\mathrm{d}s$
The result follows from (3.9) and upon observing that $g(T,x(T-))=g(t_{N},x_{N-1})$. ∎

### v2

@@ 2.1. Distribution of reaction times
With each reaction $\ell\in{\mathcal{R}}$, we associate a reaction time $\tau_{\ell}$. The reaction times satisfy, see e.g. Gillespie, (1992),
@@
If $\lambda_{\ell}$ is constant in time, it follows from (2.4) that $\tau_{\ell}\mid X(s)=x\sim\mathrm{Exp}\left(\lambda_{\ell}(x)\right)$. This implies that the time $\tau=\min_{\ell\in{\mathcal{R}}}\tau_{\ell}$ that the first reaction occurs satisfies $\tau\mid X(s)=x\sim\mathrm{Exp}\left(\sum_{\ell\in{\mathcal{R}}}\lambda_{\ell}(x)\right)$.
If the reaction times are inhomogeneous in time and 2.1 is satisfied, similarly the time for the first reaction to occur has distribution (2.4) with intensity function $\sum_{\ell\in{\mathcal{R}}}\lambda_{\ell}$. Moreover, the probability distribution of the first reaction in any subset $R\subseteq{\mathcal{R}}$ also satisfies (2.4) with intensity function $\sum_{\ell\in R}\lambda_{\ell}$.
@@ 2.2. Chemical master equation and chemical Langevin equation
with initial condition $p(s,x;s,y)=1\{y=x\}$. It is well-known, see e.g. Li, (2020) that the chemical reaction process can be approximated by solutions to the Chemical Langevin Equation (CLE), which is the SDE
@@
where (assuming the reactions to be numbered $1,\ldots,{B}$)
$\begin{split}b_{\mathrm{CLE}}(t,x)&=\sum_{\ell=1}^{B}\lambda_{\ell}(t,x)\xi_{\ell}\\ \sigma_{\mathrm{CLE}}(t,x)&=\begin{bmatrix}\xi_{1}&\ldots&\xi_{B}\end{bmatrix}\sqrt{\mbox{diag}(\lambda_{1}(t,x),\ldots,\lambda_{{B}}(t,x))}\end{split}$ (2.6)
and $W$ is an independent ${\mathbb{R}}^{B}$-valued Brownian motion.
@@ 2.3. Examples
Our simplest example models a population of initial size $x_{0}$ in which an individual dies in a time interval $(t,t+\Delta)$ with probability $c\Delta+o(\Delta)$ for some constant $c>0$ and $\Delta$ small. Such a process is modelled as chemical reaction process with just one specie and one reaction, namely $\lambda_{1}\colon(t,x)\mapsto cx$ with $\xi_{1}=-1$.
@@
Here $(x_{1},x_{2},x_{3},x_{4})$ refer to species counts of $(S,E,SE,P)$. Interesting aspects of this example are firstly that the fourth component ($P$) only appears in reaction ($3$) where $1$ is added and therefore is monotonically increasing and secondly that there are absorbing states such as $x=(0,x_{2},0,x_{4})$ where the process is killed as all reaction rates are zero. A realization of this process can be found in Figure 2.
Figure 2. Realization of the forward process of 2.4 using $\kappa_{1}=5$, $\kappa_{2}=5$, $\kappa_{3}=3$ and initial position $x_{0}=(12,10,10,10)$.
3. Guided Markov processes
3.1. First ideas
We first provide some intuition on how the dynamics of a chemical reaction network change upon conditioning on a future event. Let $\mathcal{E}$ denote some event later than time $t$, for example $\{X_{T}=v\}$. For $t\in[0,T)$, $\Delta>0$ and $x\in{\mathbb{S}}$ and $y=x+\xi_{\ell}$,
${\mathbb{P}}\left(X(t+\Delta)=y\mid X(t)=x,\mathcal{E}\right)=\frac{{\mathbb{P}}\left(X(t+\Delta)=y,\,\mathcal{E}\mid X(t)=x\right)}{{\mathbb{P}}\left(\mathcal{E}\mid X(t)=x\right)}$
$\qquad\qquad\qquad=\frac{{\mathbb{P}}\left(\mathcal{E}\mid X(t+\Delta)=y,\,X(t)=x\right){\mathbb{P}}\left(X(t+\Delta)=y\mid X(t)=x\right)}{{\mathbb{P}}\left(\mathcal{E}\mid X(t)=x\right)}$
$\qquad\qquad\qquad={\mathbb{P}}\left(X(t+\Delta)=y\mid X(t)=x\right)\frac{{\mathbb{P}}\left(\mathcal{E}\mid X(t+\Delta)=y\right)}{{\mathbb{P}}\left(\mathcal{E}\mid X(t)=x\right)}$
Taking the limit $\Delta\downarrow 0$, this suggests that if $X$ is conditioned on the event $\mathcal{E}$ then it is a chemical reaction process with adjusted intensities
$\lambda_{\ell}^{h}(t,x)=\lambda_{\ell}(t,x)\frac{h(t,x+\xi_{\ell})}{h(t,x)},$ (3.1)
where $h(t,x)={\mathbb{P}}\left(\mathcal{E}\mid X(t)=x\right)$. The process with intensities $\lambda_{\ell}^{h}(t,x)$ has generator
${\mathcal{L}}^{h}_{t}f(x)=\sum_{\ell\in\mathcal{R}}\lambda_{\ell}^{h}(t,x)\left(f(x+\xi_{\ell})-f(x)\right).$ (3.2)
To sample the conditioned process $h$ is required which is rarely available in closed form. The general idea behind the construction of what we call a guided chemical reaction process is to replace $h$ by a suitable tractable substitute $g$. This gives a process with rates $\lambda_{t}^{g}$. Discrepancies between the true conditioned process and the guided process can be accounted for by evaluating the likelihood ratio of their measures on path space. In this section we will establish sufficient conditions on $g$ for this approach to be valid.
As shown below, conditioning the process corresponds to a change of measure. To see this connection, a direct computation shows that ${\mathcal{L}}^{h}_{t}f$ as defined in (3.2) can be expressed in terms of ${\mathcal{L}}_{t}$:
$\mathcal{L}_{t}^{h}f(x)=\frac{1}{h(t,x)}\left[{\mathcal{L}}_{t}(fh)(t,x)-f(x){\mathcal{L}}_{t}h(t,x)\right].$ (3.3)
This operator is strongly connected to exponential changes of measure studied in Palmowski & Rolski, (2002) and Corstanje et al., (2023) for a wider class of Markov processes.
3.2. Guiding by a change of measure
We assume $x_{0}$ to be known and that =$\mathcal{E}=\{LX(T)=v\}$ is observed for some known $v\in{\mathbb{R}}^{m}$ and a matrix $L\in{\mathbb{R}}^{m\times d}$ of full row rank with $m\leq d$. In case $m=d$, we assume without loss of generality that $L=I$. Additionally, we assume $\mathbb{P}(LX(T)=v\mid X(0)=x_{0})>0$. The extension to multiple observations will follow in a straightforward way in later sections.
For a suitable $g\colon[0,T)\times{\mathbb{S}}\to{\mathbb{R}}$ define
$E^{g}(t)=\frac{g(t,X(t))}{g(0,x_{0})}\exp\left(-\int_{0}^{t}\frac{{\mathcal{A}}g}{g}(s,X(s))\,\mathrm{d}s\right),\qquad t\in[0,T).$ (3.4)
@@ 3.2. Guiding by a change of measure
We write $g\in{\mathcal{G}}$ if $g\in{\mathcal{D}}({\mathcal{A}})$ is a strictly positive function and there exists a filtration $({\mathcal{F}}_{t})_{t}$ such that $E^{g}(t),\,t\in[0,T)$ is a martingale adapted to $({\mathcal{F}}_{t})_{t}$.
As $({\mathcal{F}}_{t}^{X})_{t}=\sigma(X(s):s\leq t)$ is a sub-$\sigma$-algebra of $({\mathcal{F}}_{t})_{t}$, $E^{g}(t)$ is a martingale with respect to $({\mathcal{F}}_{t}^{X})_{t}$ as well whenever $g\in{\mathcal{G}}$.
Let $g\in{\mathcal{G}}$ and let ${\mathbb{P}}_{t}$ denote the law of the process $X$ restricted to ${\mathcal{F}}_{t}^{X}$. Define the new collection of measures $\{{\mathbb{P}}_{t}^{g}:t\in[0,T)\}$ on $(\Omega,{\mathcal{F}}_{t}^{X})$ by
$\,\mathrm{d}{\mathbb{P}}_{t}^{g}=E^{g}(t)\,\mathrm{d}{\mathbb{P}}_{t},\qquad t\in[0,T).$ (3.5)
By Theorem 4.2 of Palmowski & Rolski, (2002), $f(t,X(t))-\int_{0}^{t}\left(\partial_{t}f+{\mathcal{L}}^{g}_{t}f\right)(s,X(s))\,\mathrm{d}s$ is a martingale under ${\mathbb{P}}^{g}_{t}$ for any function $f(t,x)$ finitely supported in $x$ and differentiable in $t$. Hence, under $\mathbb{P}^{g}_{t}$, $X$ is a chemical reaction process with the same jumps as under ${\mathbb{P}}_{t}$ but with adjusted intensities $\lambda^{g}_{\ell}(t,x)=\lambda(t,x)g(t,x+\xi_{\ell})/g(t,x)$. For the reader’s convenience we have summarised some of the main arguments of Palmowski & Rolski, (2002) for establishing this connection in Appendix D.
@@
Some authors refer to this process as the twisted process (see e.g. Del Moral & Penev, (2017)). Upon denoting the transition probabilities of $X$ by $p$, i.e. ${\mathbb{P}}(X_{T}\in A\mid X_{t}=x)=\sum_{y\in A}p(t,x;T,y)$ for $A\subseteq{\mathbb{S}}$, it follows from Example 2.4 of Corstanje et al., (2023) that the conditioned process $\left(X\mid LX(T)=v\right)$ is obtained by taking $g=h$ with
$h(t,x)=\sum_{\zeta\in L^{-1}v}p(t,x;T,\zeta),$ (3.6)
where $L^{-1}v=\{y\in{\mathbb{S}}:Ly=v\}$. We recall our assumption $h(0,x_{0})>0$, as otherwise conditioned paths have probability zero. Note that $h$ is bounded and satisfies Kolmogorov’s backward equation: ${\mathcal{A}}h=0$. By Proposition 3.2 in Palmowski & Rolski, (2002), $h\in{\mathcal{G}}$ and thus we can define the probability measure ${\mathbb{P}}^{h}_{t}$ by
$\,\mathrm{d}{\mathbb{P}}_{t}^{h}=\frac{h(t,X(t))}{h(0,x_{0})}\,\mathrm{d}{\mathbb{P}}_{t},\qquad{t\leq T}.$
Intuitively, ${\mathbb{P}}_{t}^{h}$ gives more mass to paths where $h(t,X(t))$ is large.
Unfortunately, $h$ is intractable as the transition probabilities $p$ are only known in closed-form in very specific cases.
To resolve this, consider another function $g\in{\mathcal{G}}$ that acts as a tractable substitute for $h$. Then, as a consequence, the process $X$ is tractable and can be simulated under the measure ${\mathbb{P}}_{t}^{g}$. Moreover, for $t<T$,
$\frac{\,\mathrm{d}{\mathbb{P}}^{h}_{t}}{\,\mathrm{d}{\mathbb{P}}^{g}_{t}}(X)=\frac{E_{t}^{h}}{E_{t}^{g}}=\frac{h(t,X(t))}{g(t,X(t))}\frac{g(0,x_{0})}{h(0,x_{0})}\Psi_{t}^{g}(X),$ (3.7)
where
$\Psi_{t}^{g}(X)=\exp\left(\int_{0}^{t}\frac{{\mathcal{A}}g}{g}(s,X(s))\,\mathrm{d}s\right).$ (3.8)
To evaluate $\Psi^{g}$, a direct computation yields
Equation 3.7 becomes particularly useful if it can be evaluated in $t=T$ as well. Since $h(T,x)=\mathbf{1}_{\{Lx=v\}}$ is defined in $T$ this only depends on the choice for $g$. We present a class of functions that can be used for this purpose in 3.4.
Lemma 3.3.
Suppose $g\in{\mathcal{D}}({\mathcal{A}})$ is a strictly positive function such that for some positive constant $C$
$\int_{0}^{T}\sum_{\ell\in{\mathcal{R}}}\lambda_{\ell}(s,X(s))\left(\frac{g(s,X(s)+\xi_{\ell})}{g(s,X(s))}-1\right)^{2}\,\mathrm{d}s<C,$
$\mathbb{P}$-almost surely, then $g\in{\mathcal{G}}$.
Proof.
The proof is inspired by the proof given in Example 15.2.10 in Brémaud, (2020) and Lemma 19.6 in Liptser & Shiryaev, (2013).
It suffices to show that ${\mathbb{E}}\left[E^{g}(T)\right]=1$. By Lemma 3.1 of Palmowski & Rolski, (2002), $E^{g}$ is a local martingale. Let $\left\{\sigma_{n}\right\}_{n}$ be a localizing sequence for $E^{g}$. Then, if $(E^{g}(t)\,t\in[0,T])$ were uniformly integrable, then
${\mathbb{E}}\left[E^{g}(T)\right]=\lim_{n\to\infty}{\mathbb{E}}\left[E^{g}(T\wedge\sigma_{n})\right]={\mathbb{E}}\left[E^{g}(0)\right]=1$
We proceed to show uniform integrability. Given $t\in[0,T]$, suppose a trajectory $X$ has jumps at times $t_{1},\dots,t_{N}$. Set $t_{0}=0$ and $t_{N+1}=t$. Then
$E^{g}(t)$ $=\frac{g(t,X(t))}{g(0,x_{0})}\exp\left(-\sum_{j=0}^{N}\int_{t_{j}}^{t_{j+1}}\partial_{s}\log g(s,X(t_{j}))\,\mathrm{d}s-\int_{0}^{t}\frac{{\mathcal{L}}g(s,X(s))}{g(s,X(s))}\,\mathrm{d}s\right)$ (3.10)
$=\frac{g(t,X(t))}{g(0,x_{0})}\prod_{j=0}^{N}\frac{g(t_{j},X(t_{j}))}{g(t_{j+1},X(t_{j}))}\exp\left(-\int_{0}^{t}\frac{{\mathcal{L}}g(s,X(s))}{g(s,X(s))}\,\mathrm{d}s\right)$
$=\prod_{j=0}^{N-1}\frac{g(t_{j+1},X(t_{j+1}))}{g(t_{j+1},X(t_{j}))}\exp\left(-\int_{0}^{t}\frac{{\mathcal{L}}g(s,X(s))}{g(s,X(s))}\,\mathrm{d}s\right)$
$=\prod_{j=0}^{N-1}\frac{g(t_{j+1},X(t_{j+1}))}{g(t_{j+1},X(t_{j}))}\exp\left(-\int_{0}^{t}\sum_{\ell\in{\mathcal{R}}}\lambda_{\ell}(s,X(s))\left(\frac{g(s,X(s)+\xi_{\ell})}{g(s,X(s))}-1\right)\,\mathrm{d}s\right)$
From (3.10), it is easy to verify that
$E^{g}(t)^{2}=E^{g^{2}}(t)\exp\left(\int_{0}^{t}\sum_{\ell\in{\mathcal{R}}}\lambda_{\ell}(s,X(s))\left(\frac{g(s,X(s)+\xi_{\ell})}{g(s,X(s))}-1\right)^{2}\,\mathrm{d}s\right)$
By Lemma 3.1 of Palmowski & Rolski, (2002), $E^{g^{2}}$ is a local martingale and, since it is bounded from below, it is a supermartingale with ${\mathbb{E}}\left[E^{g^{2}}(t)\right]\leq 1$. Hence, for all $n$,
$\mathbb{E}\left[E^{g}(T\wedge\sigma_{n})^{2}\right]$
$\quad\leq{\mathbb{E}}\left[E^{g^{2}}(T\wedge\sigma_{n})\exp\left(\int_{0}^{T}\sum_{\ell\in{\mathcal{R}}}\lambda_{\ell}(s,X(s))\left(\frac{g(s,X(s)+\xi_{\ell})}{g(s,X(s))}-1\right)^{2}\,\mathrm{d}s\right)\right]\leq e^{C}$
Therefore $\sup_{n}\mathbb{E}\left[E^{g}(T\wedge\sigma_{n})^{2}\right]<\infty$. ∎
Theorem 3.4.
Define $h$ is as in (3.6) and let $g\colon[0,T]\times{\mathbb{S}}\to{\mathbb{R}}$ be such that the condition of 3.3 is satisfied. Then
$\frac{\,\mathrm{d}{\mathbb{P}}^{h}_{T}}{\,\mathrm{d}{\mathbb{P}}^{g}_{T}}=\frac{g(0,x_{0})}{h(0,x_{0})}\frac{\Psi_{T}^{g}(X)}{g(T,X(T)}\mathbf{1}_{\{LX(T)=v\}}.$ (3.11)
Proof.
The form of the Radon-Nikodym derivative follows from (3.7) upon noting that for $t\uparrow T$,
$h(t,X(t))={\mathbb{E}}\left[\mathbf{1}\{LX(T)=v\}\mid{\mathcal{F}}_{t}\right]\to{\mathbb{E}}\left[\mathbf{1}\{LX(T)=v\}\mid{\mathcal{F}}_{T}\right]=\mathbf{1}\{LX(T)=v\}.$
The following proposition is often helpful in establishing that the condition of 3.3 is satisfied for a given $g$.
Proposition 3.5.
Suppose $g\colon[0,T]\times{\mathbb{S}}\to{\mathbb{R}}$ is such that for all $x$ the map $t\mapsto g(t,x)$ is bounded from above and bounded away from zero. Then the condition of 3.3 is satisfied.
Proof.
Since we assume throughout $X$ to be nonexplosive, we have that almost surely the maps $t\mapsto g(t,X(t)+\xi_{\ell})/g(t,X(t)$ are almost surely bounded on $[0,T]$ for all $\ell\in{\mathcal{R}}$. The result now follows upon noting that $\int_{0}^{T}\sum_{\ell\in{\mathcal{R}}}\lambda_{\ell}(s,X(s))\,\mathrm{d}s<\infty$ by Assumption (2.1c). ∎
Corollary 3.6.
The likelihood of $x_{0}$ based on the observation $v$ is given by
$h(0,x_{0})=g(0,x_{0}){\mathbb{E}}_{T}^{g}\left[\frac{\Psi_{T}^{{g}}(X)}{g(T,X(T)}\mathbf{1}_{\{LX(T)=v\}}\right].$ (3.12)
Proof.
Note that $g\equiv 1$ satisfies the condition in 3.3. This choice simply yields the original forward process. Sampling the conditioned process in this way is however very inefficient when ${\mathbb{P}}\left(LX(T)=v\right)$ is low.
The lemma below shows that the law of the guided process does not change when $g$ is multiplied by a function only depending on time.
@@
Suppose $g\in{\mathcal{G}}$ and let $c\colon[0,T]\to{\mathbb{R}}_{+}$ be a differentiable function. Then $E^{cg}(t)=E^{g}(t)$.
@@
From this, we get $E^{cg}(t)$ by negating the right-hand-side, taking the exponent and subsequently multiplying by $c(t)g(t,X(t))/(c(0)g(0,x_{0}))$. It is easily seen that the terms with $c$ cancel out. ∎
The following proposition is useful for numerical evaluation of the likelihood of a sampled guided process.
@@
$\begin{split}&\frac{g(0,x_{0})}{g(T,x(T))}\exp\left(\int_{0}^{T}\frac{{\mathcal{A}}g}{g}(s,x(s))\,\mathrm{d}s\right)\\ &\qquad=\left(\prod_{j=0}^{N-2}\frac{1}{\alpha_{j}(t_{j+1})}\right)\exp\left(\sum_{j=0}^{N-1}\int_{t_{j}}^{t_{j+1}}\sum_{\ell\in{\mathcal{R}}}\lambda_{\ell}(s,x_{j})\left[\alpha_{j}(s)-1\right]\,\mathrm{d}s\right)\end{split}$
@@
This is a consequence of (3.10). ∎